# Avoid Catastrophic Forgetting with Rank-1 Fisher from Diffusion Models

### v1

@@ 3.4 Promoting parameter sharing across tasks via generative distillation
To mitigate this limitation and strengthen our evaluation, we promote parameter sharing by adding a generative distillation term computed on replayed inputs from earlier tasks (Masip et al., 2025). Concretely, we keep a frozen teacher model $\varepsilon_{\theta_{T-1}^{\star}}$ from the previous task and sample replay inputs $\tilde{x}$ from it. Generative distillation then encourages the current model $\varepsilon_{\theta}$ to match the teacher model’s denoising behavior:
@@ 4.1 Datasets for class-incremental continual learning
In class-incremental continual learning, a dataset is partitioned into $T$ tasks, where each task $T_{k}$ contains $n$ class labels and $\cap T_{k}=\emptyset$. We evaluate our approach on four image datasets commonly used in generative modeling: MNIST (Lecun et al., 1998), Fashion MNIST (FMNIST) (Xiao et al., 2017), CIFAR-10 (Krizhevsky, 2009), and ImageNet-1k (Deng et al., 2009). We use the down-sampled ImageNet-1k (Chrabaszcz et al., 2017) such that each image has the dimension of $3\times 32\times 32$ for faster training and evaluation while preserving similar performance characteristics to the full-size ImageNet-1k. We additionally pad MNIST and FMNIST to $32\times 32$ to be consistent with CIFAR-10 and down-sampled ImageNet-1k. For each of MNIST, FMNIST, and CIFAR-10, we partition the dataset into 5 tasks with 2 classes per task. For ImageNet-1k, we partition it into 20 tasks with 50 classes per task, simulating a much longer horizon continual learning task. We use the same class label ordering as in the original dataset. We refer to Appendix D.3 for additional details.
@@ 4.3 Implementation details
We use the label conditioning UNet implementation from Huggingface with default hyper-parameters as our denoising network. We use 4 ResNet blocks with $128$ output channels in the first down-sample block and $256$ output channels in the rest. For sampling, we use a DDIM scheduler with 50 sampling steps and 1000 noising steps. We train our UNet with 100 epochs on ImageNet-1k and 200 epochs on the rest of the datasets for each task with a batch size of 128 and a learning rate of $2\times 10^{-4}$ using Adam (Kingma & Ba, 2017). We use an EWC penalty weight of $15000$ for all tasks. For generative distillation, we create a replay buffer of 1300 images per class for ImageNet-1k and 5000 images per class for other datasets. We refer to Appendix C for additional training details.
@@ 5.1 Continual learning performance
Using generative distillation alone substantially improves both average FID and forgetting on all datasets compared to EWC-only by encouraging the model to move toward an optima that performs well across all (including replayed) tasks. When combined with our rank-1 EWC, we observe further improvements on all datasets. On MNIST and FMNIST, catastrophic forgetting is nearly eliminated with forgetting $\mathcal{F}=0.6_{\pm 0.1}$ and $\mathcal{F}=0.9_{\pm 0.3}$, respectively. On the long horizon ImageNet-1k dataset, rank-1 EWC halves forgetting relative to generative distillation alone ($\mathcal{F}=15.2_{\pm 0.1}$ vs. $\mathcal{F}=31.6_{\pm 0.2}$). And across all datasets, image generation quality improves substantially, further narrowing the gap to the non-continual upper bound. In contrast, while combining diagonal EWC with distillation improves the diagonal EWC-only approach, the improvement stems largely from distillation, and performance often matches or even degrades compared to distillation alone, suggesting the ineffectiveness of the diagonal EWC constraint. Overall, these results demonstrate that rank-1 EWC effectively complements generative distillation by enforcing parameter constraints within the shared optimum that distillation provides.
@@
In addition, on the long horizon ImageNet-1k dataset, both generative distillation-only and the diagonal variant begin to diverge around task 10 while the average FID only gradually increases for our rank-1 approach. In particular, the generative distillation-only approach reached a plateau around task 6 to 11 before diverging. This suggests that distillation suffers from distribution shift due to errors that have accumulated from the previous imperfect replay samples, such that the model optima moves away from early tasks. We refer Section˜D.4 for additional results.
@@ 6 Conclusion, limitations, and future work
Assumption˜1 suggests that the collinearity of per-sample gradients is dependent on the denoising network’s architecture. While current work focuses on the UNet-based diffusion models, as this is a popular architecture, future work should expand our analysis to more recent architectures, such as those used in Transformer-based diffusion models (Peebles & Xie, 2023).
@@
We used open-source datasets and model implementations as described in Section˜3.2 and Section˜4, and additional experiment details, including empirical validation experiment implementations, class-incremental dataset splits, and hyper-parameters in Appendices˜C and D.3.
@@ B.2 Proof for proposition 2
Let $g(\theta;x_{t})=\nabla_{\theta}\mathcal{L}_{\text{DSM}}(\theta;x_{t})$, $\mu(\theta)=\mathbb{E}_{x_{t}^{\prime}\sim q_{t}}\big[g(\theta;x_{t}^{\prime})\big]$. By Proposition˜1, $s^{\star}(x_{t})\approx x_{t}\gamma_{t}$, where $\gamma_{t}=-1/(1-\bar{\alpha_{t}})$ is a scalar independent of $x_{t}$. Near convergence, the denoising network approximates the true score function: $s_{\theta}(x_{t},t)\approx s^{\star}(x_{t})\approx x_{t}\gamma_{t}$. By Assumption˜1, $s_{\theta}(x_{t},t)$ takes a linear form as $s_{\theta}(x_{t},t)=x_{t}\theta$. Substitute back to $g(\theta;x_{t}):$
@@
$=\nabla_{\theta}\left(\frac{1-\bar{\alpha_{t}}}{2}\big\|x_{t}\theta-x_{t}\gamma_{t}\big\|^{2}_{2}\right)$ (9)
$=(1-\bar{\alpha_{t}})\big(x_{t}\theta-x_{t}\gamma_{t}I\big)x_{t}^{\top}$ (10)
$=(1-\bar{\alpha_{t}})\|x_{t}\|^{2}(\theta-\gamma_{t}I).$ (11)
Let $c(x_{t})=(1-\bar{\alpha_{t}})\|x_{t}\|^{2}$ be a scalar function dependent on $x_{t}$, and $v=\theta-\gamma_{t}I$ be a vector in parameter space doesn’t depend on $x_{t}$. Then $g(\theta;x_{t})=c(x_{t})v$. By the definition of collinearity, $g(\theta;x_{t})$ collinear with $v$ for any $x_{t}\sim q_{t}$. And since per-sample gradient collinear with each other, it collinear with the population mean $\mu(\theta)$. ∎
Appendix C Empirical Results
@@
Training is performed with a batch size of $128$ and learning rate $2\times 10^{-4}$ using the Adam optimizer (Kingma & Ba, 2017). Each task is trained for $200$ epochs on the full dataset. On MNIST (Lecun et al., 1998), training the base model required approximately $97$ minutes on a single Nvidia A100 (40GB) GPU.
C.1 Additional Model Variants
@@
Appendix D Continual Learning Results
D.1 Additional Implementation Details
We use the label-conditioned UNet implementation from Huggingface as our denoising network. The UNet employs three ResNet blocks with $16$ output channels in each downsampling block. For training, we use a batch size of $128$ and a learning rate of $2\times 10^{-4}$ with the Adam optimizer (Kingma & Ba, 2017), and train for 200 epochs per task for each dataset.
@@ D.1 Additional Implementation Details
D.2 Additional implementation details for continual learning experiments
@@
Optimizer Adam (Kingma & Ba, 2017)
@@
D.3 Additional Dataset Details
@@
D.4 Continual Learning Curves for each Task and Dataset

### v2

@@ 3.4 Promoting parameter sharing across tasks via generative distillation
To mitigate this limitation and strengthen our evaluation, we promote parameter sharing by adding a generative distillation term computed on replayed inputs from earlier tasks (Masip and others, 2025). Concretely, we keep a frozen teacher model $\varepsilon_{\theta_{T-1}^{\star}}$ from the previous task and sample replay inputs $\tilde{x}$ from it. Generative distillation then encourages the current model $\varepsilon_{\theta}$ to match the teacher model’s denoising behavior:
@@ 4.1 Datasets for class-incremental continual learning
In class-incremental continual learning, a dataset is partitioned into $T$ tasks, where each task $T_{k}$ contains $n$ class labels and $\cap T_{k}=\emptyset$. We evaluate our approach on four image datasets commonly used in generative modeling: MNIST (Lecun et al., 1998), Fashion MNIST (FMNIST) (Xiao et al., 2017), CIFAR-10 (Krizhevsky, 2009), and ImageNet-1k (Deng et al., 2009). We use the down-sampled ImageNet-1k (Chrabaszcz et al., 2017) such that each image has the dimension of $3\times 32\times 32$ for faster training and evaluation while preserving similar performance characteristics to the full-size ImageNet-1k. We additionally pad MNIST and FMNIST to $32\times 32$ to be consistent with CIFAR-10 and down-sampled ImageNet-1k. For each of MNIST, FMNIST, and CIFAR-10, we partition the dataset into 5 tasks with 2 classes per task. For ImageNet-1k, we partition it into 20 tasks with 50 classes per task, simulating a much longer horizon continual learning task. We use the same class label ordering as in the original dataset. We refer to Appendix E.3 for additional details.
@@ 4.3 Implementation details
We use the label conditioning UNet implementation from Huggingface with default hyper-parameters as our denoising network. We use 4 ResNet blocks with $128$ output channels in the first down-sample block and $256$ output channels in the rest. For sampling, we use a DDIM scheduler with 50 sampling steps and 1000 noising steps. We train our UNet with 100 epochs on ImageNet-1k and 200 epochs on the rest of the datasets for each task with a batch size of 128 and a learning rate of $2\times 10^{-4}$ using Adam (Kingma and Ba, 2017). We use an EWC penalty weight of $15000$ for all tasks. For generative distillation, we create a replay buffer of 1300 images per class for ImageNet-1k and 5000 images per class for other datasets. We refer to Appendix D for additional training details.
@@ 5.1 Continual learning performance
Using generative distillation alone substantially improves both average FID and forgetting on all datasets compared to EWC-only by encouraging the model to move toward an optima that performs well across all (including replayed) tasks. When combined with our rank-1 EWC, we observe further improvements on all datasets. On MNIST and FMNIST, catastrophic forgetting is nearly eliminated with forgetting $\mathcal{F}=0.6_{\pm 0.1}$ and $\mathcal{F}=0.9_{\pm 0.3}$, respectively. On the long horizon ImageNet-1k dataset, rank-1 EWC more than halves forgetting relative to generative distillation alone ($\mathcal{F}=15.2_{\pm 4.8}$ vs. $\mathcal{F}=46.2_{\pm 12.9}$). And across all datasets, image generation quality improves substantially, further narrowing the gap to the non-continual upper bound. In contrast, while combining diagonal EWC with distillation improves the diagonal EWC-only approach, the improvement stems largely from distillation, and performance often matches or even degrades compared to distillation alone, suggesting the ineffectiveness of the diagonal EWC constraint. Overall, these results demonstrate that rank-1 EWC effectively complements generative distillation by enforcing parameter constraints within the shared optimum that distillation provides. We provide additional analysis of low-rank Fisher estimation (ranks 1155) via power iteration, as well as continual fine-tuning baselines, in Section˜E.4.
@@
In addition, on the long horizon ImageNet-1k dataset, both generative distillation-only and the diagonal variant begin to diverge around task 10 while the average FID only gradually increases for our rank-1 approach. In particular, the generative distillation-only approach reached a plateau around task 6 to 11 before diverging. This suggests that distillation suffers from distribution shift due to errors that have accumulated from the previous imperfect replay samples, such that the model optima moves away from early tasks. We refer Section˜E.5 for additional results.
@@ 6 Conclusion, limitations, and future work
While the main contribution of our work lies in analyzing the gradient geometry of the diffusion model’s mathematical framework, an important direction for future work is to conduct a more in-depth empirical and theoretical study of Assumption˜1 across multiple model architectures. Such an analysis would help clarify which architectural components (e.g., skip connections, attention mechanisms, etc.) or training regularizations (e.g., L2 norms) make Assumption˜1 more or less likely to hold.
@@
We used open-source datasets and model implementations as described in Section˜3.2 and Section˜4, and additional experiment details, including empirical validation experiment implementations, class-incremental dataset splits, and hyper-parameters in Appendices˜D and E.3.
@@ B.2 Proof for proposition 2
Let $g(\theta;x_{t})=\nabla_{\theta}\mathcal{L}_{\text{DSM}}(\theta;x_{t})$, $\mu(\theta)=\mathbb{E}_{x_{t}^{\prime}\sim q_{t}}\big[g(\theta;x_{t}^{\prime})\big]$. By Proposition˜1, $s^{\star}(x_{t})\approx x_{t}\gamma_{t}$, where $\gamma_{t}=-1/(1-\bar{\alpha_{t}})$ is a scalar independent of $x_{t}$. Near convergence, the denoising network approximates the true score function: $s_{\theta}(x_{t},t)\approx s^{\star}(x_{t})\approx x_{t}\gamma_{t}$. By Assumption˜1, $s_{\theta}(x_{t},t)$ takes a linear form as $s_{\theta}(x_{t},t)=A_{\theta}x_{t}$. Substitute back to $g(\theta;x_{t}):$
@@
$=\nabla_{\theta}\left(\frac{1-\bar{\alpha_{t}}}{2}\big\|A_{\theta}x_{t}-x_{t}\gamma_{t}\big\|^{2}_{2}\right)$ (9)
$=(1-\bar{\alpha_{t}})\big(A_{\theta}x_{t}-x_{t}\gamma_{t}I\big)x_{t}^{\top}$ (10)
$=(1-\bar{\alpha_{t}})\|x_{t}\|^{2}(A_{\theta}-\gamma_{t}I).$ (11)
Let $c(x_{t})=(1-\bar{\alpha_{t}})\|x_{t}\|^{2}$ be a scalar function dependent on $x_{t}$, and $v=A_{\theta}-\gamma_{t}I$ be a vector in parameter space doesn’t depend on $x_{t}$. Then $g(\theta;x_{t})=c(x_{t})v$. By the definition of collinearity, $g(\theta;x_{t})$ collinear with $v$ for any $x_{t}\sim q_{t}$. And since per-sample gradient collinear with each other, it collinear with the population mean $\mu(\theta)$. ∎
Appendix C Additional discussion on Assumption 1
The theoretical rationale behind Assumption˜1 comes from the analysis of autoencoders. Let $f_{\theta}$ be an autoencoder trained with the standard MSE objective to approximate the identity mapping, i.e., $f_{\theta}(x)\approx x$. As stated in  Proposition˜1, in the late diffusion steps the denoising network behaves approximately like such an autoencoder $f_{\theta}$.
Vincent et al. (2010) report that an autoencoder with MSE is doing PCA-like operation when there is no nonlinearity. When nonlinearities are introduced, they argue that the learned representation is still “likely” to lie in a PCA subspace. For alternative reconstruction losses such as cross-entropy, Vincent et al. (2010) note that the autoencoder is no longer the same as that of PCA and will likely learn different features. In our setting, this suggests that the observed approximately rank-1 behavior is primarily a consequence of the specific DDPM parameterization: a non-parameterized, variance-preserving forward process that reduces the ELBO to a simple MSE objective. In particular, this rank-1 behavior need not hold when the training objective is not purely MSE.
To illustrate this, we also experimented with a VAE (Kingma and Welling, 2014) trained with an $\mathrm{MSE}+\mathrm{KL}$ objective. To quantify rank-1 dominance, we measure the percentage of total variance explained by the largest eigenvalue of the empirical Fisher information matrix:
$\frac{\lambda_{1}}{\sum_{i}\lambda_{i}}.$
When analyzing the gradients of this VAE, we found that the top eigenvalue accounts for only about $50\%$ of the total variance. In contrast, for DDPM in late timesteps, the top eigenvalue explains roughly $99\%$ of the variance.
Moreover, when we down-weight the KL term by a factor of $10^{-3}$, the top eigenvalue in the VAE setting explains about $85\%$ of the variance, suggesting that the near rank-1 behavior is closely tied to the autoencoding regime with an MSE-dominated loss.
This is consistent with our theoretical analysis: even if $f_{\theta}(x)\approx A_{\theta}x$ as it works in a PCA subspace for some affine parameter $A_{\theta}$, the gradient of a combined $\mathrm{MSE}+\mathrm{KL}$ objective cannot in general be factored as $\nabla_{\theta}\mathcal{L}(x)=c(x)\,v$, where $c(x)$ is a scalar depending on $x$ and $v$ is a parameter-space vector independent of $x$. Such a factorization is precisely the condition for gradient collinearity established in Appendix B.2.
Although a rigorous mathematical proof of this behavior for large-scale nonlinear autoencoder architectures is currently out of reach, we assume that the skip connections in the U-Net (Ronneberger et al., 2015) make it even more likely that the large-scale network effectively operates in such a PCA-like subspace (Ghorbani et al., 2019). Intuitively, these skip connections allow information to bypass some nonlinear activations, thereby preserving approximately linear structure in the mapping, in line with the observations of Vincent et al. (2010).
Appendix D Additional Empirical Results
@@
Training is performed with a batch size of $128$ and learning rate $2\times 10^{-4}$ using the Adam optimizer (Kingma and Ba, 2017). Each task is trained for $200$ epochs on the full dataset. On MNIST (Lecun et al., 1998), training the base model required approximately $97$ minutes on a single Nvidia A100 (40GB) GPU.
D.1 Additional Model Variants
@@
Appendix E Additional Continual Learning Results
E.1 Additional Implementation Details
We use the label-conditioned UNet implementation from Huggingface as our denoising network. The UNet employs three ResNet blocks with $16$ output channels in each downsampling block. For training, we use a batch size of $128$ and a learning rate of $2\times 10^{-4}$ with the Adam optimizer (Kingma and Ba, 2017), and train for 200 epochs per task for each dataset.
@@ E.1 Additional Implementation Details
E.2 Additional implementation details for continual learning experiments
@@
Optimizer Adam (Kingma and Ba, 2017)
@@
E.3 Additional Dataset Details
@@
E.4 Additional Baselines
E.4.1 Ablations on low-rank Fisher surrogates
A key practical benefit of our analysis is that it yields an efficient and scalable estimate of the leading Fisher eigen-direction using only the empirical mean gradient $\mathbb{E}[g]$, avoiding explicit formation of the Fisher and the costly iterative procedures typically required to extract top eigenvectors in high-dimensional parameter spaces. To contextualize this design choice, we compare against low-rank baselines that estimate the top-$k$ Fisher subspace via stochastic power iteration and then construct rank-$k$ approximations (rank-1 through rank-5). Across both CIFAR-10 and FashionMNIST, these iterative low-rank baselines do not improve performance over our mean-gradient rank-1 estimate, and in many cases perform worse, consistent with the difficulty of reliably converging power iteration under limited compute. Moreover, increasing the approximation rank beyond 1 yields little benefit, supporting the view that the curvature is already well-captured by the dominant direction in this regime.
Table 5: Low-rank ablations using rank-$k$ Fisher approximations (estimated via stochastic power iteration) versus our mean-gradient rank-1 estimate. We report final FID and average forgetting (mean $\pm$ standard error).
CIFAR-10 FashionMNIST
Method Final FID Avg Forgetting Final FID Avg Forgetting
GD + rank-1 (ours) $50.5\pm 1.2$ $7.4\pm 1.2$ $15.4\pm 0.6$ $0.9\pm 0.3$
GD + rank-2 $57.5\pm 2.6$ $15.9\pm 1.8$ $18.1\pm 1.6$ $3.8\pm 1.4$
GD + rank-3 $60.4\pm 5.4$ $17.7\pm 5.0$ $16.8\pm 0.9$ $2.9\pm 0.6$
GD + rank-4 $55.7\pm 3.4$ $13.0\pm 2.0$ $17.0\pm 1.2$ $2.9\pm 1.2$
GD + rank-5 $55.6\pm 2.3$ $12.6\pm 1.6$ $17.1\pm 1.2$ $2.9\pm 1.0$
E.4.2 Continual fine-tuning without regularization
We additionally report the performance of plain continual fine-tuning with no regularization (i.e., setting the penalty weight to $\lambda=0$). This baseline isolates the extent of catastrophic forgetting in the absence of any stabilization mechanism, and serves as a sanity check that our regularized baselines are not benefiting from overly conservative hyperparameters. As shown in Table˜6, removing regularization yields substantially worse final generation quality and severe forgetting across datasets.
Table 6: Continual fine-tuning with no regularization ($\lambda=0$). We report final FID and average forgetting.
Dataset Final FID Avg Forgetting
CIFAR-10 115.410 72.496
FashionMNIST 103.955 88.009
ImageNet32 265.902 67.864
MNIST 102.642 93.700
E.5 Continual Learning Curves for each Task and Dataset